# Unified model for non-Abelian braiding of Majorana and Dirac fermion zero modes

### v1

@@ Abstract
One can notice that for case III in TABLE 1 that $\varphi_{2}=\varphi_{3}=0$, the cross-coupling terms like $i\gamma_{i}\widetilde{\gamma}_{j}$ vanish. In this way, a unitary symmetry $\mathcal{R}$ defined as $\mathcal{R}_{ij}i\gamma_{i}\widetilde{\gamma}_{j}\mathcal{R}_{ij}^{-1}=-i\gamma_{i}\widetilde{\gamma}_{j}$ is preserved Hong et al. (2022), and all these MZMs are separated into two sets as shown in Fig. 3(a). According to the braiding protocol introduced above, these two independent sets of MZMs $\{\gamma_{i}\}$ and $\{\widetilde{\gamma}_{i}\}$ are braided simultaneously. It is worth noticing that $\gamma_{i}$ and $\widetilde{\gamma}_{i}$ are spatially fully-overlapped. Therefore, the low-lying modes being braided here are actually fully-overlapped Majorana pairs respecting local unitary symmetry $\mathcal{R}$. Then each Majorana pair as a whole is actually Dirac fermion zero mode Hong et al. (2022); Wu et al. (2020a, 2023) that the corresponding braiding properties can be constructed from the Majorana ones.
In the adiabatic limit $T\gg\frac{1}{w},\frac{1}{\Delta}$ [$w$ is the (average) energy difference between the lowest two-fold states and the second lowest two-fold states, see Fig. 3(c)], the Berry phase for the lowest two-fold states turns out to be the projection of the tensor product of the Majorana one’s into the parity-even subspace as ^$\hat{\mathbb{P}}_{\mathrm{even}}[\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}]=\mathrm{diag}\{-i,i\}$. Similarly, the Berry phase for the second lowest two-fold states is ^$\hat{\mathbb{P}}_{\mathrm{even}}[\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes\mathrm{diag}\{e^{i\pi/4},e^{-i\pi/4}\}]=\mathrm{diag}\{1,1\}$ Sup . In addition to this, the dynamic phases also contribute when $T\gg\frac{1}{w},\frac{1}{\Delta}$, and hence the braiding operator reads $U_{\mathrm{case~{}III}}=e^{i\theta}\left[e^{-i\delta}\begin{pmatrix}-i&0\\ 0&i\end{pmatrix}\oplus e^{i\delta}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right]$. Here, $\theta$ ($\delta$) are the dynamic phases corresponding to the average energy of (energy difference between) the lowest and second lowest two-fold dengenerate states.
By contrast, for $1/w\gg T\gg 1/\Delta$, the dynamic phase of the four lowest states only contributes a trivial overall phase and can be dropped. Moreover, the Berry phase for the Majorana set $\{\widetilde{\gamma}_{j}\}$ now reduces to an identity $I_{4\times 4}$ since $\widetilde{\gamma}_{j}$’s are braided in the non-adiabatic limit $1/w\gg T$ Chang (2021). Therefore, in this case, the braiding operator is obtained by projecting the tensor product $\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes I_{4\times 4}$ into the parity-even subspace, giving $U_{\mathrm{case~{}III}}=\mathrm{diag}\{e^{-i\pi/4},e^{-i\pi/4},e^{i\pi/4},e^{i\pi/4}\}$ Sup .
@@
Finally, for case IV in TABLE 1 that $\varphi_{2},\varphi_{3}\neq 0$, the unitary symmetry $\mathcal{R}$ is broken and the two sets of MZMs $\{\gamma_{i}\}$ and $\{\widetilde{\gamma}_{i}\}$ are coupled via the cross-coupling terms [see Fig. 3(b)]. The energy spectra in Fig. 3(d) shows that all the two-fold degeneracies are still preserved, while the energies of these degenerate states vary during the braiding. Through numerical calculation Sup , we have confirmed that the Berry phase accumulated in case IV is exactly the same as the one in case III (up to a gauge transform). This is in analogy with the consequence stated above that the Berry phase is the same for case II and case I. The difference lies in that now the dynamic phases $\theta\mp\delta$ are $\varphi_{2}$-, $\varphi_{3}$-, and $|\Omega_{j}|/|\chi_{j}|$-dependent since the energy during the braiding vary with $\varphi_{2}$, $\varphi_{3}$, and $|\Omega_{j}|/|\chi_{j}|$. Meanwhile, the braiding results still exhibit an oscillation with increasing $T$ Sup . Moreover, the (average) energy difference $w$ between the two-fold ground states and the second-lowest two-fold states also increases with $\varphi_{2}$, $\varphi_{3}$, and $\big{|}|\chi_{j}|-|\Omega_{j}|\big{|}$. In this way, the $1/w\gg T\gg 1/\Delta$ condition can no longer be satisfied when $\varphi_{2}$, $\varphi_{3}$, or $\big{|}|\chi_{j}|-|\Omega_{j}|\big{|}$ is significantly non-zero. As a result, an additional dynamic phase is always accumulated alongside the geometric phase, providing a clear signature of the symmetry-breaking effect Wu et al. (2020a, 2023, b) on the braiding of Dirac fermion zero modes.
@@
Discussions. The minimal Kitaev chain model is experimentally feasible, in ultracold atom systems for instance, especially for $|\Omega_{j}|=0\neq|\chi_{j}|$, as the complexity of pairing term is circumvented. In this case, the low-lying modes are Dirac fermion modes with unitary symmetry preserved (case III) or broken (case IV), other than MZMs. In this condition, the Hamiltonian in Eq. (1) can be engineered by utilizing the internal degree of freedom of atoms. Taking 171Yb fermions as an example, the four braiding modes correspond to the atomic spin states, 1S0$|F=1/2,m_{F}=\pm 1/2\rangle$, and ${}^{3}P$0$|F=1/2,m_{F}=\pm 1/2\rangle$, which have significantly long lifetime. The tunnelings among the four fermion modes are highly controllable by manipulating Raman and clock transitions Jenkins et al. (2022).
@@
Acknowledgements. This work was financially supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302400 and No. 2021ZD0302000), the National Natural Science Foundation of China (Grant No. 12304194, No. 12425401 and No. 12261160368), National Key Research and Development Program of China (2021YFA1400900), and by Shanghai Municipal Science and Technology (Grant No. 24DP2600100).

### v2

@@ Abstract
One can notice that for case III in TABLE 1 that $\varphi_{2}=\varphi_{3}=0$, the cross-coupling terms like $i\gamma_{i}\widetilde{\gamma}_{j}$ vanish. In this way, a unitary symmetry $\mathcal{R}$ defined as $\mathcal{R}_{ij}i\gamma_{i}\widetilde{\gamma}_{j}\mathcal{R}_{ij}^{-1}=-i\gamma_{i}\widetilde{\gamma}_{j}$ is preserved Hong et al. (2022), and all these MZMs are separated into two sets as shown in Fig. 3(a). According to the braiding protocol introduced above, these two independent sets of MZMs $\{\gamma_{i}\}$ and $\{\widetilde{\gamma}_{i}\}$ are braided simultaneously. It is worth noticing that $\gamma_{i}$ and $\widetilde{\gamma}_{i}$ are spatially fully-overlapped. Therefore, the low-lying modes being braided here are actually fully-overlapped Majorana pairs respecting local unitary symmetry $\mathcal{R}$. Then each Majorana pair as a whole is actually complex fermion mode Wu et al. (2023) that the corresponding braiding properties can be constructed from the Majorana ones.
In the adiabatic limit $T\gg 1/w,1/\Delta$ [$w$ is the (average) energy difference between the lowest two-fold states and the second lowest two-fold states, see Fig. 3(c)], the Berry phase for the lowest two-fold states turns out to be the projection of the tensor product of the Majorana one’s into the parity-even subspace as ^$\hat{\mathbb{P}}_{\mathrm{even}}[\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}]=\mathrm{diag}\{-i,i\}$. Similarly, the Berry phase for the second lowest two-fold states is ^$\hat{\mathbb{P}}_{\mathrm{even}}[\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes\mathrm{diag}\{e^{i\pi/4},e^{-i\pi/4}\}]=\mathrm{diag}\{1,1\}$ Sup . Meanwhile, the dynamic phases also contribute. Hence, the braiding operator reads $U_{\mathrm{case~{}III}}=e^{i\theta}\left[e^{-i\delta}\begin{pmatrix}-i&0\\ 0&i\end{pmatrix}\oplus e^{i\delta}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right]$, where $\theta$ ($\delta$) are the dynamic phases corresponding to the average energy of (energy difference between) the lowest and second lowest two-fold degenerate states.
By contrast, for $1/w\gg T\gg 1/\Delta$, the dynamic phase of the four lowest states only contributes a trivial overall phase. Moreover, the Berry phase for the Majorana set $\{\widetilde{\gamma}_{j}\}$ is an identity $I_{4\times 4}$ here since $\widetilde{\gamma}_{j}$’s are braided in the non-adiabatic limit $1/w\gg T$ Chang (2021). Therefore, in this case, the braiding operator is obtained by projecting $\mathrm{diag}\{e^{-i\pi/4},e^{i\pi/4}\}\otimes I_{4\times 4}$ into the parity-even subspace, giving $U_{\mathrm{case~{}III}}=\mathrm{diag}\{e^{-i\pi/4},e^{-i\pi/4},e^{i\pi/4},e^{i\pi/4}\}$ Sup .
@@
Finally, for case IV in TABLE 1 that $\varphi_{2},\varphi_{3}\neq 0$, the unitary symmetry $\mathcal{R}$ is broken and the two sets of MZMs $\{\gamma_{i}\}$ and $\{\widetilde{\gamma}_{i}\}$ are coupled via the cross-coupling terms [see Fig. 3(b)]. The energy spectra in Fig. 3(d) shows that all the two-fold degeneracies are still preserved, while the energies of these degenerate states vary during the braiding. Through numerical calculation Sup , we have confirmed that the Berry phase accumulated in case IV is exactly the same as the one in case III (up to a gauge transform). This is in analogy with the consequence stated above that the Berry phase is the same for case II and case I. The difference lies in that now the dynamic phases $\theta\mp\delta$ are $\varphi_{2}$-, $\varphi_{3}$-, and $|\Omega_{j}|/|\chi_{j}|$-dependent since the energy spectra vary with $\varphi_{2}$, $\varphi_{3}$, and $|\Omega_{j}|/|\chi_{j}|$. Meanwhile, the braiding results still exhibit an oscillation with increasing $T$ Sup . Moreover, the (average) energy difference $w$ between the two-fold ground states and the second-lowest two-fold states also increases with $\varphi_{2}$, $\varphi_{3}$, and $\big{|}|\chi_{j}|-|\Omega_{j}|\big{|}$. In this way, the $1/w\gg T\gg 1/\Delta$ condition can no longer be satisfied when $\varphi_{2}$, $\varphi_{3}$, or $\big{|}|\chi_{j}|-|\Omega_{j}|\big{|}$ is significantly non-zero. Consequently, an additional dynamic phase is always accumulated alongside the geometric phase, a clear signature of the symmetry-breaking effect Wu et al. (2020b) on the braiding of Dirac fermion zero modes.
@@
Discussions. The “sweet spot” condition has long been a key objective in experiments Bordin et al. (2024); ten Haaf et al. (2024); Bordin et al. (2025); Liu et al. (2022). We have shown here the non-trivial phases with symmetry protected MZM pairs or Dirac zero modes in the regime away from this condition. With the removal of fine-tuning condition $|\Omega_{j}|=|\chi_{j}|$, the minimal Kitaev chain model becomes more realistic for experimental realization. Particularly, our prediction is of high experimental feasibility by setting $|\chi_{j}|>|\Omega_{j}|=0$, as the complexity associated with the superconducting pairing is circumvented. In this case, the low-lying modes being braided are Dirac fermion modes that the fermion number is conserved during the braiding. The minimal Kitaev chain has recently been experimentally engineered in QD arrays Bordin et al. (2024); ten Haaf et al. (2024); Bordin et al. (2025) and could also be implemented by utilizing the internal degrees of freedom of ultracold atoms. Taking 171Yb fermions as an example, the four braiding modes correspond to the atomic spin states, 1S0$|F=1/2,m_{F}=\pm 1/2\rangle$, and ${}^{3}P$0$|F=1/2,m_{F}=\pm 1/2\rangle$, which have significantly long lifetime. The tunnelings among the four fermion modes are highly controllable and readily realized by manipulating Raman and clock transitions Jenkins et al. (2022). This study paves the way for the ultimate realization of non-Abelian statistics in experiment.
@@
Acknowledgements. The authors thank Leo P. Kouwenhoven and Chun-Xiao Liu for fruitful discussions. This work was financially supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302400 and No. 2021ZD0302000), the National Natural Science Foundation of China (Grant No. 12304194, No. 12425401 and No. 12261160368), National Key Research and Development Program of China (Grant No. 2021YFA1400900 and No. 2024YFA1409000), Shanghai Municipal Science and Technology (Grant No. 24DP2600100), Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01), and by Shanghai Science and Technology Innovation Action Plan (Grant No. 24LZ1400800).